RC Meeting and Runplan, Monday, Oct 1
- The accelerator did not get beam back until around 5pm Sunday. Most of the work was getting good transmission in the injector.
- Finished the elastic singles at around 6pm
- While switching to ep coincidence (SHMS change polarity) took optics data in the HMS ( Run 4826).
- Start coincidence ep data taking. Had missing reference time errors in replay. Brad came in to investigate. It seems that "a slightly loose lemo end on the hEL-real delay loop feeding the T5 coin. input." cause the problem. See https://logbooks.jlab.org/entry/3601696 for details. This highlights the importance of monitoring the replay for report of missing reference times.
- In late swing and early owl, accelerator fixed problems with 5 pass separator and other work. beam around 50uA after that.
- Finished ep coincidence. Took about 25K coincidence ep elastics.
- Around 10:30 switched to first physics run. Finished first LH2 data point at around 4. https://logbooks.jlab.org/entry/3602123
- No RF studies on Tuesday. Hopefully next will be next Tuesday. About 6 RF trips/hr.
- BCM calibration done. Put it in gscaler.param . https://logbooks.jlab.org/entry/3602122
- Accelerator plans to slowly increase current from 50 to 65uA.
- Have about 40 hours of physics taking before aerogel change.
- Mike Fowler fixed PID loop for the HMS Dipole JT3 file.
- Plan to do the aerogel change on Friday.